# iSeg: An Iterative Refinement-based Framework for Training-free Segmentation

### v2

@@ 3.3 Our iterative refinement framework
Here we introduce our proposed iterative refinement framework for training-free segmentation, named iSeg. We employ pre-trained stable diffusion model for feature extraction, and perform segmentation using the generated self-attention and cross-attention maps with iterative refinement. To address performance degradation issue in iterative refinement, we introduce an entropy-reduced self-attention (Ent-Self) module to suppress irrelevant global information within self-attention map. In addition, we design a category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention map as the input of iterative refinement. The algorithm is detailed in Appendix A.
@@
where represents the denoising U-net, and indicates the timestep that is set as a fixed value like [16]. Then, we can extract self-attention and cross-attention maps at different layers of denoising UNet. Similar to [44], we only select the high-resolution self-attention map at last decoder layer, and upsample and fuse the cross-attention maps at different decoder layers as initial cross-attention map . For self-attention map , we introduce an entropy-reduced self-attention (Ent-Self) module to reduce irrelevant global information within original self-attention map, where the entropy-reduced self-attention map is represent as . Then, we perform the iterative refinement using entropy-reduced self-attention map as
where represents the total number of iterations, and is the original cross-attention map . After each iteration, we add a normalization operation. For simplicity, we do not show the normalization in Eq. 3 and Fig. 3. Finally, we can extract the segmentation map of a category at the index of text prompts as . It is worth noting that an accurate initial cross-attention map is also important, which provides a better input for iterative refinement. Therefore, instead of original cross-attention map, we design a simple category-enhanced cross-attention (Cat-Cross) module to obtain more accurate cross-attention map fed to iterative refinement.
Entropy-reduced self-attention module. As mentioned above, the weak responses of irrelevant regions within self-attention map contain the global information and interfere with the refined cross-attention map. To address this issue, we introduce an entropy-reduced self-attention (Ent-Self) module. Following [54], we calculate self-attention entropy by viewing the attention distribution as a probability mass function of a discrete random variable, which can be written as
@@
where represents the updating factor. The entropy-reduced operation makes the value of near to one becomes larger, and makes the value of near to zero becomes smaller. As a result, the entropy-reduced operation can suppress the weak responses within self-attention map corresponding to irrelevant global information, and does not reduce the strong responses corresponding to the local region.
@@ 4 Experiments
We perform the experiments on weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation for synthetic dataset. Following the prior works [15, 16], we mainly use pixel accuracy (ACC) and mean intersection over union (mIoU) to evaluate segmentation performance.
| Method | Publication | Training | VOC | COCO | |
@@
| baseline (Ours) | - | 66.0 | 39.4 | ||
| Diffusion-based | iSeg (Ours) | - | 73.6 | 44.7 |
Implementation details. We employ the pre-trained stable diffusion V2.1 as the base model, and feed the input images into the pre-trained model for segmentation without any training on segmentation dataset. The input images are resized to 512512 pixels. All segmentation tasks are evaluated on a single NVIDIA RTX 3090 GPU with 24G memory. During inference, only 5G memory is required when using half precision with a batch size of 1, resulting in an approximate inference time of 0.16 seconds per image. For weakly-supervised semantic segmentation, we evaluate pseudo mask generation and use the categories of corresponding training images to generate text prompts. For open-vocabulary semantic segmentation, we employ pre-trained CLIP to generate text prompts. For unsupervised segmentation and mask generation for synthetic dataset, we employ our entropy-reduced self-attention module to refine the segmentation results of DiffSeg [15] and DiffuMask [18]. Please refer to Appendix B for more implementation details of different segmentation tasks.
| Method | Publication | Training | VOC | Context | Object | |
@@
| OVDiff [46] | arXiv2023 | 67.1 | 30.1 | 34.8 | ||
@@
| baseline (Ours) | - | 61.4 | 26.4 | 36.1 | ||
| Diffusion-based | iSeg (Ours) | - | 66.0 | 30.6 | 38.2 |
@@ 4.1 State-of-the-art comparison
Weakly supervised semantic segmentation. Table 1 compares our training-free method iSeg and some weakly-supervised methods, including CNN-based, Transformer-based, CLIP-based, and Diffusion-based approaches, on both PASCAL VOC [68] and MS COCO [69] training sets. Here, we only compare the quality of generated pseudo masks. Our proposed iSeg achieves the mIoU scores of 73.6% and 44.7% on PASCAL VOC and MS COCO, respectively. Compared to diffusion-based training-free DiffSegmenter [16] and T2M [17], our iSeg has 3.1% and 0.9% improvements in terms of mIoU on PASCAL VOC. Our proposed iSeg also outperforms the training-based approaches. For instance, compared to ToCo [61] and WeakTr [62], our proposed iSeg has 1.4% and 7.4% improvements in terms of mIoU on PASCAL VOC.
@@
Open-vocabulary semantic segmentation. Table 2 compares our proposed method with some open-vocabulary semantic segmentation methods, including CLIP-based and diffusion-based approaches, on PASCAL VOC validation set, PASCAL Context validation set [70], and MS COCO Object validation set. Our proposed iSeg outperforms the training-free DiffSegmenter [16] by an absolute gain of , , and on these three datasets.
| Method | Training | Cityscapes | COCO-Stuff-27 | ||
| ACC | mIoU | ACC | mIoU | ||
| IIC [39] | 47.9 | 6.4 | 21.8 | 6.7 | |
| MDC [71] | 40.7 | 7.1 | 32.3 | 9.8 | |
| PICLE [72] | 65.5 | 12.3 | 48.1 | 13.8 | |
| STEGO [40] | 73.2 | 21.0 | 56.9 | 28.2 | |
| MaskCLIP [63] | 35.9 | 10.0 | 32.2 | 19.6 | |
| RoCo [41] | 74.6 | 19.3 | 46.1 | 26.3 | |
| DiffSeg [15] | 76.0 | 21.2 | 72.5 | 43.6 | |
| iSeg (Ours) | 78.7 | 25.0 | 74.5 | 45.2 | |
@@
Unsupervised segmentation. Table 3 compares our method with some unsupervised segmentation methods on Cityscapes [73] validation set and COCO-Stuff-27 [74] validation set. We employ our iterative refinement as a post-processing, and integrate it into unsupervised DiffSeg [15] to improve segmentation results. Our proposed iSeg can also present stable improvement on both Cityscapes and COCO-Stuff-27. For instance, our iSeg outperforms DiffSeg by a large gain of in terms of ACC and in terms of mIoU on Cityscapes.
@@
Mask generation on synthetic dataset. Fig. 5 presents some qualitative results of our proposed iSeg and DiffuMask [18]. The synthetic images are from DiffuMask [18]. DiffuMask first generates coarse pseudo masks using cross-attention map, and then train an AffinityNet to generate more precise pseudo masks. Compared to DiffuMask, our proposed iSeg has more accurate segmentation results without using any training processing. More results on synthetic images can be found in appendix C.
@@ 4.2 Ablation study
| Ent-Self | Cat-Cross | mIoU |
| 66.0 | ||
| 70.1 | ||
| 68.9 | ||
| 73.6 |
Impact of integrating two proposed modules. Table LABEL:ablation:modules shows the results on PASCAL VOC training set. When integrating entropy-reduced self-attention (Ent-Self) module, it achieves a performance improvement of 4.1%. When integrating category-enhanced cross-attention (Cat-Cross) module, it presents a performance improvement of 2.9%. By integrating both Ent-Self and Cat-Cross into the baseline, it totally improves the baseline by 7.6%, which demonstrates that two proposed modules are complementary. Specifically, Cat-Cross module provides a better input for refinement, while Ent-Self module avoids the performance degradation during refinement.
@@
Hyper-parameter settings. There are some hyper-parameters in our method, including the updating factor in our Ent-Self module, the number of iterations , and the weighting factor in our Cat-Cross module. Fig. 6(a) presents the impact of and . When (namely using original self-attention map), iterative refinement presents a fast performance drop after several iterations. With a relatively large (e.g., 0.05), iterative refinement initially improves performance, and then keeps relatively stable performance. When and , it achieves best performance. Fig. 6(b) presents the impact of weighting factor . Increasing initially leads to improved performance but eventually results in performance drop. The best performance is achieved when .
4.3 Segmentation with different kinds of images and interactions
Fig. 7 shows that our method supports segmentation on different kinds of images and interactions. At left, our method performs accurate segmentation of given categories on natural, synthetic, clipart, painting and sketch images. At right, our method can support different interactions to perform segmentation, where the initial segmentation map is generated by unsupervised segmentation maps, point, box, and line, and then fed to our iterative refinement with Self-Ent module. More results are in Appendix C. We also provide a video demo in supplementary material.

### v3

@@ 3.3 Our iterative refinement framework
Here we introduce our proposed iterative refinement framework for training-free segmentation, named iSeg. We employ pre-trained stable diffusion model for feature extraction, and perform segmentation using the generated self-attention and cross-attention maps with iterative refinement. To address performance degradation issue in iterative refinement, we introduce an entropy-reduced self-attention (Ent-Self) module to suppress irrelevant global information within self-attention map. We find that, our iterative refinement with entropy-reduced module can be also viewed as to generate the refined self-attention map closer to optimal solution (i.e., the ground-truth self-attention map). In addition, we design a category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention map as the input of iterative refinement.
@@
where represents the denoising UNet, and indicates the timestep that is set as a fixed value like [63]. Then, we can extract self-attention and cross-attention maps at different layers of denoising UNet. Similar to [27], we only select the high-resolution self-attention map at last decoder layer, and upsample and fuse the cross-attention maps at different decoder layers as initial cross-attention map . For self-attention map , we introduce an entropy-reduced self-attention (Ent-Self) module to reduce irrelevant global information within original self-attention map, where the entropy-reduced self-attention map is represent as . Then, we perform iterative refinement using entropy-reduced self-attention map as
where represents the total number of iterations. is the original cross-attention map corresponding to category names, which is selected from initial cross-attention map according to the category indices. After each iteration, we add a normalization operation. For simplicity, we do not show the normalization in Eq. 4 and Fig. 3. Finally, we generate the segmentation map by binarizing the refined cross-attention map .
It is worth noting that an accurate initial cross-attention map is also important, which provides a better input for iterative refinement. Therefore, instead of original cross-attention map, we design a simple category-enhanced cross-attention (Cat-Cross) module to obtain more accurate cross-attention map fed to iterative refinement.
Entropy-reduced self-attention module. As mentioned above, the weak responses of irrelevant regions within self-attention map contain the global information and interfere with the refined cross-attention map. To address this issue, we introduce an entropy-reduced self-attention (Ent-Self) module. Following [25], we calculate self-attention entropy by viewing the attention distribution as a probability mass function of a discrete random variable, which is written as
@@
where represents the updating factor.
Analysis on effectiveness. The reason why our iterative refinement with entropy-reduced self-attention module leads to better segmentation results can be explained from two perspectives. (i) The entropy-reduced operation makes the value of near to one becomes larger, and makes the value of near to zero becomes smaller. As a result, the entropy-reduced operation can suppress the weak responses within self-attention map corresponding to irrelevant global information, and does not reduce the strong responses corresponding to the local region. With this entropy-reduced operation, our iterative refinement can progressively enhance the local object region within cross-attention map without highlighting irrelevant regions during multiple iterations. (ii) Our iterative refinement can be viewed as to generate the better refined self-attention map, which is then used to refine cross-attention map once. Specifically, our iterative refinement can be approximately written as , where can be seen as a refined self-attention map after iterations. We observe that the refined self-attention map becomes clean and closer to the ground-truth self-attention map after multiple iterations. Fig. 4 shows some examples of self-attention map at selected points. Therefore, compared to original self-attention map, the refined self-attention map can better improve cross-attention map.
@@
Algorithm 1 provides the pseudo code about how to refine the cross-attention by entropy-reduced self-attention iteratively using our proposed iSeg.
@@ 4 Experiments
In this section, to validate the effectiveness and superiority of our method, we perform the experiments on weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation for synthetic dataset. Following the prior works [59, 63], we mainly use pixel accuracy (ACC) and mean intersection over union (mIoU) to evaluate segmentation performance.
Type | Method | Publication | Training | VOC | COCO |
@@
Diffusion-based | iSeg (Ours) | - | 75.2 | 45.5 |
4.1 Implementation Details
We employ the pre-trained stable diffusion V2.1 as the base model, and feed the input image into the pre-trained model for feature extraction without any training on segmentation dataset. The input image is resized to 512512 pixels. All segmentation tasks are evaluated on a single NVIDIA RTX 3090 GPU with 24G memory. During inference, only 5G memory is required when using half precision with a batch size of 1, resulting in an approximate inference time of 0.16 seconds per image. We present the implementation details on various segmentation tasks below.
Weakly-supervised semantic segmentation. In this task, one key is to generate pseudo pixel-level masks for training set. We compare our training-free method with existing approaches for pixel-level mask generation on two public datasets, including PASCAL VOC 2012 [18] and MS COCO 2014 [36]. Specifically, we use the categories existing in the image to generate a text template such as ‘A photograph of [CLS A] and [CLS B] and other objects and background’ and add the background categories, similar to CLIPES [37].
Type | Method | Publication | Training | VOC | Context | Object |
@@ 4.1 Implementation Details
CaR [58] | CVPR2024 | 67.6 | 30.5 | 36.6 | ||
SAM-based | SAM-CLIP [61] | CVPRW2024 | 60.6 | 29.2 | 31.5 | |
OVDiff [26] | ECCV2024 | 67.1 | 30.1 | 34.8 | ||
@@
Diffusion-based | iSeg (Ours) | - | 68.2 | 30.9 | 38.4 |
Open-vocabulary semantic segmentation. This task aims to perform segmentation of arbitrary categories. We compare our training-free method with existing approaches on three public datasets, including PASCAL VOC 2012 (VOC), PASCAL-Context (Context), and MS COCO Object (Object). To perform open-vocabulary semantic segmentation, we employ the pre-trained CLIP to generate image-level category labels as text prompts following TagCLIP [38]. Afterwards, we keep the similar settings as weakly-supervised semantic segmentation task to obtain the final segmentation masks of different categories.
Unsupervised segmentation. This task groups pixels into different objects without using any annotations. Here we compare our method with other unsupervised segmentation approaches on two public datasets, including COCO-Stuff-27 [5] and Cityscapes [15]. We integrate our proposed method into the training-free approach DiffSeg [59] as a post-processing. Specifically, we treat the segmentation results of DiffSeg as initial cross-attention maps, and employ our proposed entropy-reduced self-attention module to perform iterative refinement, while the other settings are consistent with DiffSeg.
Mask generation for synthetic dataset. Recently, researchers have explored generating synthetic dataset for semantic segmentation. In this task, it is important to generate precise mask labels for synthetic images. DiffuMask [65] utilizes cross-attention maps in diffusion model to generate pseudo masks for synthetic images, which are further refined by AffinityNet. Here, we can directly apply our iterative refinement to improve the initial pseudo masks generated by DiffuMask on synthetic dataset without requiring any additional refinements like AffinityNet.
Method | Publication | Training | Cityscapes | COCO-Stuff-27 | ||
ACC | mIoU | ACC | mIoU | |||
MDC [7] | ECCV2018 | 40.7 | 7.1 | 32.3 | 9.8 | |
IIC [23] | ICCV2019 | 47.9 | 6.4 | 21.8 | 6.7 | |
PICLE [14] | CVPR2021 | 65.5 | 12.3 | 48.1 | 13.8 | |
STEGO [42] | ICLR2022 | 73.2 | 21.0 | 56.9 | 28.2 | |
MaskCLIP [79] | ECCV2022 | 35.9 | 10.0 | 32.2 | 19.6 | |
RoCo [57] | NeurIPS2022 | 74.6 | 19.3 | 46.1 | 26.3 | |
ACSeg [33] | CVPR2023 | - | - | - | 28.1 | |
DiffSeg [59] | CVPR2024 | 76.0 | 21.2 | 72.5 | 43.6 | |
iSeg (Ours) | - | 78.7 | 25.0 | 74.5 | 45.2 |
@@ 4.2 State-of-the-art comparison
Weakly supervised semantic segmentation. Table I compares our proposed training-free method iSeg and some weakly-supervised methods, including CNN-based, transformer-based, CLIP-based, and diffusion-based approaches, on both PASCAL VOC [18] and MS COCO [36] training sets. Here, we compare the quality of generated pseudo masks of different methods on training set. Our proposed iSeg achieves the mIoU scores of 75.2% and 45.5% on PASCAL VOC and MS COCO, respectively. Compared to diffusion-based training-free methods DiffSegmenter [63] and T2M [67], our iSeg has 4.7% and 2.5% improvements in terms of mIoU on PASCAL VOC. Compared to CLIP-based method CLIP-ES [37], our proposed method iSeg has 4.4% and 5.8% improvements in terms of mIoU on PASCAL VOC and MS COCO. In addition, our proposed iSeg outperforms some training-based approaches. For instance, compared to training-based methods ToCo [54] and WeakTr [80], our proposed iSeg has 3.0% and 9.0% improvements in terms of mIoU on PASCAL VOC.
@@
Open-vocabulary semantic segmentation. Table II compares our proposed method iSeg with some open-vocabulary semantic segmentation methods, including CLIP-based, SAM-based and diffusion-based approaches, on PASCAL VOC validation set, PASCAL Context validation set [43], and MS COCO Object validation set. Our proposed iSeg achieves the mIoU scores of 68.2%, 30.9% and 38.4% on these three datasets, respectively. Compared to training-free DiffSegmenter [63], our proposed iSeg has the absolute gain of , , and on these three datasets, respectively. Compared to training-free OVDiff [26], our proposed iSeg has the absolute gain of , , and , respectively. In addition, our diffusion-based iSeg outperforms CLIP-based approaches, such as SegCLIP [41], ViewCo [51], and TagCLIP [38]. For instance, our iSeg outperforms ViewCo [51] by the absolute gain of , , and on these three datasets. The CLIP-based CaR [58] achieves comparable performance with our proposed method iSeg. However, CaR suffers from slow inference speed (1s per image) due to heavy feature extraction at multiple iterations. Moreover, our training-free iSeg is superior to training-based SAM-CLIP [61], which outperforms SAM-CLIP [51] by the absolute gain of , , and on three datasets, respectively.
Ent-Self | Cat-Cross | Weakly-supervised | Open-vocabulary | Unsupervised | ||||
VOC | COCO | VOC | Context | Object | Cityscapes | COCO-Stuff | ||
68.2 | 40.1 | 63.7 | 26.4 | 36.6 | 22.8 | 44.4 | ||
72.0 | 42.5 | 67.1 | 28.2 | 37.5 | 25.0 | 45.2 | ||
75.2 | 45.5 | 68.2 | 30.9 | 38.4 | N/A | N/A |
@@
Unsupervised segmentation. Table III compares our method with some unsupervised segmentation methods on Cityscapes [15] validation set and COCO-Stuff-27 [5] validation set. We employ our iterative refinement as a post-processing, and integrate it into unsupervised DiffSeg [59] to improve segmentation results. Our proposed iSeg presents stable improvement on both Cityscapes and COCO-Stuff-27 validation sets. For instance, our iSeg outperforms DiffSeg by a large gain of in terms of ACC and in terms of mIoU on Cityscapes. In addition, our proposed method outperforms some training-based methods. For instance, our iSeg outperforms MaskCLIP [79] and RoCo [57] by and 5.7% in terms of ACC on Cityscapes and COCO-Stuff-27, respectively.
@@
Mask generation on synthetic dataset. Fig. 6 gives some qualitative results of our proposed iSeg and DiffuMask [65]. The synthetic images (a) are from DiffuMask [65]. For mask generation, DiffuMask first generates coarse pseudo masks (b) using cross-attention map, and then trains an AffinityNet to generate more precise pseudo masks (c). Compared to using AffinityNet, our proposed iSeg has more accurate segmentation results (d) without using any training processing. For better visualization, we zoom the local segmentation results of AffinityNet and our iSeg in (e) and (f). For instance, our proposed iSeg provides more segmentation about dog outline in second row.
@@ 4.3 Ablation study
Here we perform the ablation study to demonstrate the effectiveness of our proposed modules.
Impact of integrating two proposed modules. Table LABEL:ablation:modules shows the results on various datasets and tasks. To demonstrate the effectiveness of our iterative refinement, the baseline simply refines cross-attention map with self-attention map only once. When integrating entropy-reduced self-attention (Ent-Self) module, it achieves the performance improvement on all the tasks and datasets. For instance, it presents 3.8% improvement on VOC for weakly-supervised semantic segmentation, 3.4% improvement on VOC for open-vocabulary semantic segmentation, and 2.2% improvement on CityScapes for open-vocabulary semantic segmentation. By further integrating both Ent-Self and Cat-Cross into the baseline, it has the best performance. For instance, it improves the baseline by 7.0% on VOC for weakly-supervised semantic segmentation, and 4.5% on both VOC and COCO Object for open-vocabulary semantic segmentation. It note that, our Cat-Cross module is not used for unsupervised segmentation task, because this unsupervised task does not requires category information.
@@
Hyper-parameter settings. There are some hyper-parameters in our method, including the number of iterations, the updating factor in our Ent-Self module, and the weighting factor in our Cat-Cross module. We conduct the ablation study of hyper-parameter setting on VOC for weakly-supervised semantic segmentation. Table Vc(a) presents the impact of . With the increasing iterations, it has the performance improvement. When the number , it has the best result. Table Vc(b) shows the impact of updating factor . When the updating factor , it achieves the best performance. When the updating factor becomes larger or smaller, it has some performance drop. Table Vc(c) gives the impact of weighting factor . When the weighting factor , it has the best performance.
1 | 2 | 4 | 6 | 8 | 10 | 12 | |
mIoU | 71.0 | 72.9 | 74.5 | 75.0 | 75.1 | 75.2 | 74.9 |
0 | 0.001 | 0.005 | 0.01 | 0.05 | 0.1 | |
mIoU | 69.1 | 74.3 | 75.0 | 75.2 | 74.6 | 74.1 |
1 | 1.2 | 1.4 | 1.6 | 1.8 | 2 | |
mIoU | 72.0 | 73.6 | 74.7 | 75.2 | 75.2 | 74.9 |
Level | 1616 | 3232 | Both |
mIoU | 74.8 | 56.9 | 75.2 |
Layer | #-3 | #-2 | #-1 |
mIoU | 68.5 | 71.1 | 75.2 |
Number | 1 | 50 | 100 | 150 | 200 |
mIoU | 73.2 | 74.6 | 75.2 | 74.5 | 74.3 |
Cross-attention map after different iterations. Fig. 7 visualizes some refined cross-attention maps of our proposed iSeg at different iterations. With the increment of iterations, our proposed iSeg can generate more accurate cross-attention maps of corresponding objects. For instance, after multiple iterations, our method can make cross-attention map contour more precise and uniform, which benefits the high-quality segmentation.
Impact of different designs. Table VIc further gives the impact of different designs in our method, including different layers (a,b) for cross-attention and self-attention maps, and different time-steps (c). In (a), we give the impact of using cross-attestation map at different levels. It demonstrates that, the fused cross-attention map of 1616 and 3232 pixels have the best result. In (b), we show the impact of different self-attention maps. #-1 represents self-attention map at last decoder layer. #-3 and #-2 fuse self-attention map of last decoder layer with that of third and second last decoder layer, respectively. It has the best result when using self-attention map at last decoder layer. In (c), we perform the experiments using different time-steps. We set time-step as 100 as our final settings, which has the best result.
4.4 Qualitative Results
We show segmentation results of our proposed method on natural images, synthetic images, and unsupervised segmentation task.
Results on natural images. Fig. 8 presents some segmentation results of our proposed iSeg. The images are from PASCAL VOC 2012 and MS COCO. It can be observed that our proposed method can generate accurate segmentation masks for various categories. For instance, our proposed iSeg accurately segments the tiny ball in third example, and the adjacent person and umbrella in last example.
Results on synthetic images. Fig. 9 presents some segmentation results of our proposed iSeg on synthetic images which are generated by image synthesis method DiffuMask [65]. Similar to natural images, our proposed method can accurately segment different categories for diverse synthetic images. For instance, our proposed iSeg generate accurate masks for different categories, such as potted plants, chairs, horses, birds, and cars.
Results of unsupervised segmentation. Fig. 10 gives some unsupervised segmentation results of our iSeg on COCO-Stuff-27 dataset. The first and third rows present the original images, while the second and fourth rows give the segmentation results. Our proposed method is able to accurately group the pixels into different categories.
4.5 Extensions
Here, we show that our proposed method can support the segmentation for different kinds of cross-domain images, and perform the accurate segmentation using different interactions like point, line, and box.
Different kinds of images. Fig. 11 presents some kinds of image segmentation examples on DomainNet [46]. The DomainNet includes three parts: DomainNet Sketch, DomainNet Painting, and DomainNet Clipart, which are significantly different to natural images and synthetic images. Our proposed method is able to accurately segment different kinds of objects including sketch, painting and clipart according to given categories. For instance, our method accurately segments the guitar in first sketch image, bear in fourth painting image, and cactus in last clipart image.
Different interactions. Fig. 12 presents more segmentation results using different interactions, such as points, lines, and boxes. The first row in each sub-figure shows the image and corresponding interaction. We first generate the initial cross-attention maps. Specifically, we set the pixels belonging to points and lines as foreground object, or set the pixels within boxes as foreground object. By iterative refinement, our method can refine the initial cross-attention maps, and thus accurately segment the objects corresponding to points, lines, or boxes, as in second row. For instance, our proposed method accurately segments the selected laptop computer without mistaking the orange on the laptop as the laptop.